OpenStudy (anonymous):

Solve the equation 7(2b – 5) + 3 = 10

OpenStudy (christos):

Can you separate "b" from the rest as a first step?

OpenStudy (anonymous):

....yes?

OpenStudy (christos):

So what's the equation now?

OpenStudy (anonymous):

well, it's solving for b.

OpenStudy (anonymous):

oh, I see what you mean. hold on I'll try

OpenStudy (christos):

7(2b - 5) + 3 = 10 STEP 1: 7(2b - 5) = 10 - 3 STEP 2: 7(2b - 5) = 7 STEP 3: 2b - 5 = 7/7 STEP 4: 2b - 5 = 1 STEP 5 2b = 1 + 5 STEP 6: 2b = 6 STEP 7: b = 6/2 FINAL STEP: b = 3

OpenStudy (anonymous):

Thanks!

OpenStudy (christos):

No problem. Basically that's how you separate it step by step.